You have plenty of cash to invest. You are considering an investment of $125,000 in a project which is expected to earn $14425 a year for ten years. Is it an attractive investment if your minimum expected annual rate of return is 5% (compound interest)? Calculate the expected annual rate of return (R) of the project. (6 points) NO TABLE, Full calculations, explain what happened, please have the correct answers when using an equation...
Answer #1

We need to calculate the value of annuity which gives $14425 using 5% as annual rate of interest for n=10 (10 years), and compare it with the initial investment of $125000 to see if it makes sense.

Formula for NPV of annuity = P((1-(1+r)^-n))/r)

P 14425
r 5%
n 10

Hence, the NPV of this will be $111386, which is lower than $125000. It doesn't make sense to investment in this project.

We can also calculate the annual rate of return for cash flows of -125000 in period 0 and then 14425 in period 1 to 10 by equating the NPV formula we used above to 125000

So, 14425((1-(1+r)^-10))/r = 125000 and solve for r

we get r = 2.69%